OpenStudy (anonymous):

Solve Please: 28x + 8 = 28x + 8

OpenStudy (anonymous):

0=0

OpenStudy (jiteshmeghwal9):

28x-28x=8-8 0=0

OpenStudy (anonymous):

this is the reflexive property

OpenStudy (anonymous):

does that mean all real numbers?

OpenStudy (anonymous):

yes i suppose so

OpenStudy (anonymous):

thx!

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

I know that answers, i just keep forgetting what all real numbers and stuff like that are.

OpenStudy (anonymous):

the not that. sory